A tellers supply of dimes and nickels totals $3.60. If the teller's supply of dimes numbers 1 more than 3 times the number of nickels, how many dimes are there?
Answer by scott8148(6628)

"A tellers supply of dimes and nickels totals $3.60" ___ 10d + 5n = 360 ___ 5n = 360 - 10d ___ n = 72 - 2d

"dimes numbers 1 more than 3 times the number of nickels" ___ d = 3n + 1

substituting ___ d = 3(72 - 2d) + 1 ___ d = 216 - 6d + 1 ___ 7d = 217